1. What are the vector components of the following: Round off your answer to two decimal numbers.

a. G = 10m, 1500

b. L = 10m, 900

c. A = 10m, 3600

d. D = 10m, 360 N of W


1
Expert's answer
2021-09-25T13:10:43-0400

a.

"G_x=G\cos\theta=-8.66\text{ m},\\
G_y=G\sin\theta=5\text{ m}."

b.

"L_x=L\cos\theta=0\text{ m},\\
L_y=L\sin\theta=10\text{ m}."

c.

"A_x=A\cos\theta=10\text{ m},\\
A_y=A\sin\theta=0\text{ m}."

d.

"D_x=D\cos\theta=-10\text{ m},\\
D_y=D\sin\theta=0\text{ m}."